OpenStudy (anonymous):

I think im going to be on this all night omg =( ... so the question says ... Byron is 3 years older than doug. the product of their ages is 40. how old is doug? I have an idea on how to solve it but i'm not sure if i'm right . help?

jimthompson5910 (jim_thompson5910):

what do you have so far?

OpenStudy (anonymous):

so letting d= Dougs age and letting d+3= Byrons age I got d(d+3)=40 im kinda stuck there.

jimthompson5910 (jim_thompson5910):

good start

jimthompson5910 (jim_thompson5910):

now distribute d(d+3)=40 d*d+d*3=40 d^2 + 3d = 40 and then get everything to one side d^2 + 3d = 40 d^2 + 3d - 40 = 0 what's next?

OpenStudy (anonymous):

dont I have to make a t chart?  im soo confused